OpenStudy (mrbee):

Which expression is equivalent to log2(9x3) Answers can be: log2(9)+3log2(x) log2(x)+3log2(9) 3log2(x)-log2(9) 3log2(9)-log2(x)

OpenStudy (evoker):

Is the question \[\log_{2} (9\times3)\]

OpenStudy (mathmate):

or \(\log_{2} (9x^3)\)?

OpenStudy (evoker):

I agree based on answers probably what you typed @mathmate

OpenStudy (phi):

MrBee , you should use the equation editor, because it's not clear what equation you are typing in.

OpenStudy (welshfella):

or if you type in ^ for the power 9x^3

OpenStudy (welshfella):

use the laws of logarithms log ab = log a + log b log a^b = b log a

OpenStudy (welshfella):

log2(9x^3) using the first rule = log2(9) + log2(x^3)

OpenStudy (welshfella):

now apply the above second law to log2(x^3) and you'll get the answer

OpenStudy (welshfella):

- that is the law log a^b = b log a

OpenStudy (welshfella):

in our case a = x and b = 3

OpenStudy (welshfella):

just plug those values in